OpenStudy (gabbyalicorn):

Bill measures the circumference of a circular fish pond to be 28 ft. What is the diameter of the fish pond, rounded to the nearest foot? Use 3.14 to approximate pi. So, 3.14 x 28 x 2 / 2 ? am i right on how i am doing this?

OpenStudy (tkhunny):

You already know the circumference. There should be some division in there in order to find the diameter or the radius.

OpenStudy (gabbyalicorn):

yes, after i get the answer which would be 175.84 then i would divide it by 2 which would be 87.92

OpenStudy (gabbyalicorn):

@tkhunny

OpenStudy (gabbyalicorn):

am i right?

OpenStudy (michele_laino):

hint: If C is the measure of the circumference, and D is the measure of the diameter, then we can write this: \[\Large C = 3.14 \times D\] so we have: \[\Large D = \frac{C}{{3.14}} = \frac{{28}}{{3.14}} = ...feet\]

OpenStudy (tkhunny):

You didn't listen. Why would the diameter be greater than the circumference. Circumference = 2 * Radius * pi Circumference = Diameter * pi Rethink!

OpenStudy (gabbyalicorn):

brb for a sec

OpenStudy (gabbyalicorn):

okay, back

OpenStudy (gabbyalicorn):

3.14 / 28 ?

OpenStudy (michele_laino):

no, please look at formulas of @tkhunny, or at my formulas above

OpenStudy (gabbyalicorn):

i did

OpenStudy (michele_laino):

we have: \[\Large D = \frac{C}{{3.14}} = \frac{{28}}{{3.14}} = ...feet\]

OpenStudy (gabbyalicorn):

yes,